missy Elliott invests some bonds at 8.5% simple interest and $2000 more than twice the amount in a retirement account at 10% simple interest if her total amount annual income is $1625 how much did she invest in bonds 8.5%
Answer by mananth(16946)

missy Elliott invests some bonds at 8.5% simple interest and $2000 more than twice the amount in a retirement account at 10% simple interest if her total amount annual income is $1625 how much did she invest in bonds 8.5%
missy Elliott invests some bonds at 8.5% ---------x
and $2000+2x in a retirement account at 10% simple interest
8.5%x+(2x+2000)*10%=1625
multiply by 100
8.5x+20x+20000=162500
28.5x=162500-20000
28.5x=142500
/28.5
x=142500/28.5
x=$5000
missy Elliott invests some bonds at 8.5% ---------$ 5000